OpenStudy (anonymous):

Find dy/dx. 2x^2y −3/y = 0. How do I do this one?

zepdrix (zepdrix):

\[\Large\rm 2x^2y-\frac{3}{y}=0\]Is this the correct formatting?

OpenStudy (anonymous):

yes it is :)

zepdrix (zepdrix):

\[\Large\rm 2x^2y-3y^{-1}=0\]So looks like we need to start with product rule, yes?\[\Large\rm derivative:\qquad \color{royalblue}{(2x^2)'}y+2x^2\color{royalblue}{(y)'}-\color{royalblue}{(3y^{-1})'}=0\]Blue parts are where we need to take a derivative.

OpenStudy (anonymous):

ok so 1st one derivative is : 4x 2nd one is:1 3rd one is: .... -3?

OpenStudy (shamil98):

The 2nd one is 1y'

OpenStudy (shamil98):

or just y'

zepdrix (zepdrix):

...

OpenStudy (anonymous):

Then not sure how to step forward from here.. Product rule?

OpenStudy (anonymous):

@zepdrix is the answer 4xy + 2x^2 - 3/y^2?

OpenStudy (mathmate):

Example of product rule, and implicit differentiation: \((xy^2)'=x'y^2+x(y^2)'=1*y^2+x(2y*y')=y^2+2xyy'\)

OpenStudy (anonymous):

isn't it easier to use quotient rule for 3/y? it becomes 0.y - 3.1 divided by y^2 or - 3/y^2?

OpenStudy (mathmate):

@kennybm Note the y' has to be attached after if you use the chain rule to differentiate with respect to y firrst: \((y^2)'=2yy'\)

OpenStudy (anonymous):

I'm so lost right now... I'm trying to take this course online and it hasn't explained all this stuff to me..

OpenStudy (anonymous):

I will give a medal to whoever can help me...

OpenStudy (anonymous):

and a good testimony.. Somebody, please???

OpenStudy (mathmate):

Do you follow @zepdrix 's first and second posts?

OpenStudy (anonymous):

I do and I answered the first thing to do with finding derivative but what do I do now?

OpenStudy (anonymous):

I got 4x, y and then -3

OpenStudy (mathmate):

But they are not correct, and he was probably no longer on line.

OpenStudy (mathmate):

Can we start from his second post?

OpenStudy (mathmate):

Can you evaluate \((2x^2)y\)?

OpenStudy (mathmate):

4x is not wrong, but not complete! :(

OpenStudy (anonymous):

4x^1?

OpenStudy (anonymous):

4xy

OpenStudy (mathmate):

Yes, but you have a "y" stuck to it, so you have to keep it to make 4xy, right? That's part of the product rule.

OpenStudy (mathmate):

Yes, 4xy is right! Now the second term!

OpenStudy (anonymous):

do i take derivative of 2x^2? or just the y?

OpenStudy (mathmate):

just the y!

OpenStudy (anonymous):

just y would stay y right?

OpenStudy (mathmate):

It's a "trick" question. Since you don't know much about y, the derivative of y is y' ! :)

OpenStudy (mathmate):

So the second term becomes...

OpenStudy (mathmate):

@zydic you there?

OpenStudy (anonymous):

yeah, sorry, so it is now 4xy, y and then next etrm, is it -3 or how do I do that?

OpenStudy (mathmate):

The second term is not y. 2x^2(y') Recall that in the product rule, the part that you don't differentiate will still be multiplied with the part that you differentiate.

OpenStudy (anonymous):

oooooooh

OpenStudy (anonymous):

so it would be 2x^2y

OpenStudy (mathmate):

2x^2y means you did not differentiate anything. You're supposed to differentiate y to get y'. So the result is 2x^2y'.

OpenStudy (anonymous):

oh, forgot the y' instead of y part..

OpenStudy (mathmate):

ok so far before doing the third term?

OpenStudy (anonymous):

so I would have 4xy+2x^2y'

OpenStudy (mathmate):

Yes, plus the third term.

OpenStudy (anonymous):

would i carry the - before the ( )into the equation?

OpenStudy (mathmate):

It's up to you, but the result should be the same!

OpenStudy (anonymous):

ok so would it be? -3/y^2

OpenStudy (mathmate):

Almost, only you have to include the -ve sign before the parentheses. (3/y)'=-3/y^2*y' so -(3/y)'=+3/y^2*y' Do not forget that you need to use the chain rule.

OpenStudy (mathmate):

d(K)/dx = d(K)/dy * (dy/dx)

OpenStudy (mathmate):

where dy/dx = y'

OpenStudy (anonymous):

....so i have 3/y^2*-3/y^2?

OpenStudy (mathmate):

The third term is +3/y^2*y' The y' comes from the chain rule. +3/y^2 was obtained by differentiating with respect to y, but we are differentiating with respect to x. Recall the chain rule d(K)/dx = d(K)/dy * (dy/dx) = d(K)/dy * y'

OpenStudy (anonymous):

ok, ok.. I'm still catching on here.. :)

OpenStudy (mathmate):

−(3/y)′ = -d(3/y)/dy * dy/dx = -(-3/y^2) * y' = +(3/y^2)y' Would this be better?

OpenStudy (anonymous):

So it is such a long way but I am trying. so it is all together 4xy+2x^2y'++3/y^2*y'

OpenStudy (anonymous):

ok, i see it :)

OpenStudy (mathmate):

Good, now we do not forget the right hand side: 4xy+2x^2y'+(3/y^2)y' = 0

OpenStudy (mathmate):

All you need to do now is to isolate y' (which means dy/dx) and solve for y'. Can you do that?

OpenStudy (anonymous):

oooh....

OpenStudy (anonymous):

can you really differentiate a 0?

OpenStudy (mathmate):

0 is a constant, so the derivative of a constant is...

OpenStudy (anonymous):

0

OpenStudy (mathmate):

Yep! You're doing great!

OpenStudy (anonymous):

thanks, it's alot to take in... I appreciate all your patience with me.. and I appreciate all your help!

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

Calculus is a tough subject! so, since the right hand side is 0 do I do anything more?

OpenStudy (mathmate):

Since the derivative of 0 is 0, you end up with the equation: 4xy+2x^2y'+(3/y^2)y' = 0 from which you can solve for y' (=dy/dx), which is answer requested.

OpenStudy (anonymous):

are they looking for the 0 or the whole thing written up? or do I have to take more derivatives?

OpenStudy (mathmate):

They are looking for the derivative dy/dx, which equals y'. By solving the equation for y', you will have the correct answer.

OpenStudy (anonymous):

isn't this a quadratic?

OpenStudy (mathmate):

It is more complex than a quadratic in x and y, but it is linear in y'. So the solution is relative easy to find.

OpenStudy (anonymous):

ohhh, geez, I'm not sure how to solve this then..

OpenStudy (anonymous):

I'm getting more paper out...

OpenStudy (mathmate):

4xy+2x^2y'+(3/y^2)y' = 0 2x^2y'+(3/y^2)y' = -4xy y'(2x^2+3/y^2) = -4xy Can you continue to find y'?

OpenStudy (anonymous):

.is the next part y'=-4xy/2x^2+s/y^2?

OpenStudy (anonymous):

oops that bottom s is supposed to be a 3/y^2

OpenStudy (anonymous):

can't i somehow simplify the -4xy/2x^2+3/y^2?

OpenStudy (mathmate):

You do not need to.

OpenStudy (anonymous):

so that should be the final on what they are looking for?

OpenStudy (mathmate):

The answer will contain a "mixture" of x and y (but not y').

OpenStudy (mathmate):

y'(2x^2+3/y^2) = -4xy => dy/dx= y' = -4xy/(2x^2+3/y^2)

OpenStudy (anonymous):

y' is -4xy/(2x^2+3/y^2)

OpenStudy (mathmate):

That is correct!

OpenStudy (anonymous):

oh my goodness! That was such a complicated process! i hope I can get this before my midterm next week!! I've now got a headache! ;) wow, Thank you!!!! I'm now going to give you that wonderful testimony! :)

OpenStudy (mathmate):

Rest a little before you continue! You're doing great! It always pay to understand every step, or else you get stuck with the next problem. :)

OpenStudy (mathmate):

* pays

OpenStudy (anonymous):

:)